OpenStudy (anonymous):

Prove mathematically that for any two vectors u and w lying in the xy plane of a given reference frame mag²(u+w)= mag²(u)+mag²(w)+2mag(u)mag(w)cos θ where θ is the angle between directions of p and q?

OpenStudy (unklerhaukus):

let the angle \(u\) makes with the x-axis be \(\phi\)\[\vec u=u_x\,\hat x+u_y\hat y=u(\cos\phi\,\hat x+\sin\phi\,\hat y)\] and the angle \(w\) makes with the x-axis be \(\psi\)\[\vec w=w_x\,\hat x+w_y\,\hat y=w(\cos\psi\,\hat x+\sin\psi\,\hat y)\]

OpenStudy (unklerhaukus):

\[\vec u+\vec w=u(\cos\phi\,\hat x+\sin\phi\,\hat y)+w(\cos\psi\,\hat x+\sin\psi\,\hat y)\\ |\vec u+\vec w|=\sqrt{(u(\cos\phi\,\hat x+\sin\phi\,\hat y))^2+(w(\cos\psi\,\hat x+\sin\psi\,\hat y))^2}\\ |\vec u+\vec w|^2=(u(\cos\phi\,\hat x+\sin\phi\,\hat y))^2+(w(\cos\psi\,\hat x+\sin\psi\,\hat y))^2\]

OpenStudy (unklerhaukus):

now simplify and use \[\vec x\cdot\vec x = 1\\\vec y\cdot\vec y = 1\\ \vec x\cdot\vec y = 0\\ \\ \cos^2\phi+\sin^2\phi= 1\\ \cos^2\psi+\sin^2\psi= 1\]

OpenStudy (unklerhaukus):

and \(\phi -\psi= \theta\)

OpenStudy (unklerhaukus):

do you think that will work?

OpenStudy (anonymous):

Can't you just use the distributive law for taking dot products of vectors ?

OpenStudy (anonymous):

\[\left| u+w^2 \right|=(u+w).(u+w)=\left| u \right|^2+\left| w \right|^2+2\left| u \right|\left| w \right|\cos(\theta)\] QED